Which of the molecules is/are planar?
(1)F2C=C=CF2
(2)F2B-C==C-BF2
(3)(SiH3)3N
(4)NH2-NH2

Grade:

1 Answers

Yash Sinha
13 years ago

Let us start back wards!

4) NH2-NH2 Here n is sp3 hybridised, hence it is non planar

3) (SiH3)3N Here the lone pair of N has been absorbed by the vacant d- orbital. Hence it is sp2 hybridised and planar.

2) F2B-C=C=C-BF2

BF2-C= is planar and =C-BF2 is also planar. But concentrate on central C atom!. Say it uses x axis for sigma bonds. Then it uses Py and Pz for pi bonds which are inclined at 90*. Hence the 2 triagonal planar structures of BF2C are inclined at 90*.

1)F2C=C=CF2    Same as 2nd.